# xOffense: An AI-driven autonomous penetration testing framework with offensive knowledge-enhanced LLMs and multi agent systems

### v1

@@ Abstract
Penetration testing (pentest) is essential for assessing the security of computer systems, yet conventional automated approaches using Machine Learning (ML), Deep Learning (DL) or Reinforcement Learning (RL) remain limited by simplified action spaces, high computational costs and weak reasoning across the multi-stage process of reconnaissance, vulnerability analysis and exploitation. Recent frameworks such as PentestGPT, VulnBot attempt to leverage large language models (LLMs) but face challenges of high cost, scalability and adaptability in complex workflows. This motivates the transition from monolithic models toward agentic AI systems, where multiple specialized agents collaborate to complete pentest tasks with higher efficiency and accuracy. This work introduces xOffense, an AI-driven, multi-agent penetration testing framework that shifts the process from labor-intensive, expert-driven manual efforts to fully automated, machine-executable workflows capable of scaling seamlessly with computational infrastructure. At its core, xOffense leverages a fine-tuned, mid-scale open-source LLM (Qwen3-32B) to drive reasoning and decision-making in penetration testing. The framework assigns specialized agents to reconnaissance, vulnerability scanning, and exploitation, with an orchestration layer ensuring seamless coordination across phases. Fine-tuning on Chain-of-Thought penetration testing data further enables the model to generate precise tool commands and perform consistent multi-step reasoning. We evaluate xOffense on two rigorous benchmarks: AutoPenBench and AI-Pentest-Benchmark. The results demonstrate that xOffense consistently outperforms contemporary methods, achieving a sub-task completion rate of 79.17%, decisively surpassing leading systems such as VulnBot and PentestGPT. These findings highlight the potential of domain-adapted mid-scale LLMs, when embedded within structured multi-agent orchestration, to deliver superior, cost-efficient, and reproducible solutions for autonomous penetration testing.
@@
Penetration testing , Large Language Model , Agentic AI , Multi-Agent Systems, Retrieval-Augmented Generation
@@ 1 Introduction
Penetration testing (pentest) remains one of the most effective ways to assess the real-world security posture of modern information systems. Unfortunately, the prevailing approach—manual testing conducted by small teams of human experts—cannot keep pace with today’s rapidly expanding attack surface. In 2024 alone, the National Vulnerability Database listed more than 29,000 new CVEs, a 38% year-over-year increase [1] [2]. As networks grow in scale and complexity, the gap between the appearance of new vulnerabilities and the ability of security professionals to detect and remediate them is widening. This growing imbalance underscores the urgent necessity for automated and intelligent penetration testing solutions.
@@
A second, more adaptive line of work leverages Reinforcement learning (RL). Systems like IAPTS [5] and HA-DRL [6] model penetration testing as a sequential decision-making problem in partially observable environments, enabling agents to autonomously explore and learn effective attack strategies through interaction and reward-based learning. RL agents can, in principle, discover novel attack paths, yet in practice they face two key obstacles: (i) their action space must be heavily simplified (e.g., “scan port”, “exploit CVE-xxx”), and (ii) training requires a large number of environment interactions that are expensive to obtain and seldom transfer between real networks. Consequently, even state-of-the-art RL-based pentesters achieve modest coverage and require significant engineering to integrate new tools or protocols.
@@
These limitations illustrate that purely Deep Learning (DL) or RL-based automation is insufficient for the inherently multi-phase and dynamic nature of penetration testing. To overcome this, recent works have turned toward AI agent-based paradigms, in which multiple specialized agents collaborate to emulate the workflow of human red teams. In such systems, each agent assumes a distinct role: a Reconnaissance Agent focuses on host and service discovery, a Vulnerability Analysis Agent correlates findings with CVE and CWE knowledge bases, and an Exploitation Agent generates and tests candidate payloads. This agent-oriented decomposition enables modularity, context retention across phases, and the possibility of scaling to complex attack paths that traditional ML/DL/RL pipelines cannot handle.
@@
Recent advancements in Large Language Models (LLMs) have opened new possibilities for automating penetration testing. Leveraging their strong reasoning and code generation capabilities, LLMs have been adopted in several research prototypes such as PentestGPT [7], PentestAgent [8], and VulnBot [9], where models assist or autonomously conduct reconnaissance, scanning, and exploitation. In particular, VulnBot represents a major step forward: it frames penetration testing as a collaborative workflow between specialized LLM agents guided by a Penetration Task Graph (PTG), enabling the simulation of expert-level pentesting with limited or no human intervention. Empirical results from benchmarks like AutoPenBench [10] and AI-Pentest-Benchmark [11] have validated VulnBot’s capacity to outperform other automated methods in structured testing environments.
@@
However, most of these systems rely heavily on extremely large or commercial LLMs, such as GPT-4o, LLaMA3-70B, or DeepSeek-V3. Despite their capabilities, these models present significant operational hurdles—including high resource consumption, costly API dependencies, and limited adaptability to domain-specific fine-tuning. Additionally, their general-purpose nature often leads to hallucinations, loss of context across phases, or poor command translation in complex penetration workflows. As such, there is a pressing need to explore whether smaller, fine-tuned open-source models can serve as more efficient, specialized alternatives—offering better controllability, lower cost, and targeted reasoning.
@@
This work is motivated by two core observations. First, while large LLMs have demonstrated strong potential in security domains, scale alone does not guarantee effectiveness, particularly when models are deployed in structured, multi-step tasks like penetration testing. Despite their size, large-scale LLMs still suffer from context loss across phases, generate incorrect tool usage, and require significant human supervision. Second, most current systems adopt LLMs as black-box assistants or instruction followers, without integrating deeper task-specific guidance or domain adaptation. To address these limitations, we explore an alternative paradigm: leveraging a mid-sized, domain-adapted LLM that is explicitly trained for penetration testing tasks. We present xOffense, a refined evolution of the VulnBot framework that substitutes its dependence on large, general-purpose models with a fine-tuned Qwen3-32B [12], a 32-billion-parameter open-source language model. Through dedicated training on penetration testing workflows, including vulnerability scanning, exploit crafting, and security tool interaction, xOffense achieves sharper task alignment, enhanced operational fidelity, and greater adaptability in nuanced or low-visibility environments.
@@
Beyond simply replacing the core language model, xOffense also incorporates a context-aware prompting scheme we refer to as grey-box prompting. In this setup, agents are equipped with partial system insights, such as protocol hints, observed services, or prior scan summaries, enabling them to make more informed decisions without relying on full system disclosure. This strategy preserves the operational constraints of black-box testing while offering minimal structured guidance, striking a balance between realism and agent effectiveness. By preserving VulnBot’s three-phase pipeline—reconnaissance, scanning, and exploitation—xOffense ensures compatibility with existing workflows, facilitates direct benchmarking, and provides a robust foundation for comparative evaluation.
@@
In this paper, we present the design, implementation, and evaluation of xOffense, a lightweight, domain-adaptive, and highly effective autonomous penetration testing system. Our key contributions are as follows:
@@
    An AI-driven multi-agent penetration testing system. We propose a novel agent-based framework in which specialized agents collaborate to cover all critical phases of penetration testing—reconnaissance, vulnerability analysis, and exploitation. This design emulates the workflow of human red teams, ensures modularity across tasks, and enables coherent orchestration of complex attack paths in an autonomous manner.
@@
    A domain-adapted mid-scale LLM. At the core of our system lies Qwen3-32B, a 32B-parameter open-source model fine-tuned with Chain-of-Thought penetration testing data. This adaptation empowers the model with precise multi-phase reasoning, accurate tool command generation, and strong adaptability in complex exploitation workflows.
@@
    Grey-box phase prompting. We introduce a context-aware prompting mechanism that selectively integrates environmental cues—such as observed protocols, discovered services, and prior scan outputs—into the agent reasoning process. This strategy strikes a balance between black-box and white-box testing, reducing context loss and improving continuity across phases.
@@
    Extensive empirical validation. We conduct rigorous evaluations of xOffense on AutoPenBench and AI-Pentest-Benchmark, demonstrating state-of-the-art performance in both synthetic and real-world penetration testing scenarios. The system achieves superior task and sub-task completion rates compared to prior methods, confirming the effectiveness of multi-agent orchestration and domain-adapted LLMs.
The remainder of this paper is organized as follows. Section 2 reviews prior studies on penetration testing and automation approaches, while Section 3 introduces the fundamental concepts that underpin automated systems. Section 4 describes the architecture of the proposed xOffense framework, including its fine-tuned Qwen3-32B model, grey-box prompting strategy, and multi-agent orchestration. We describe the experimental settings and benchmark datasets, evaluation metrics in Section 5. In Section 6 empirical results on two benchmarks and real-world exploitation scenarios are reported and analyzed. Section 7 discusses potential threats to validity and their implications for generalizability. Finally, Section 8 concludes the paper and outlines directions for future research.
@@ 2.1 Pre-LLM automation and RL
Deterministic orchestrators, such as DeepExploit, integrate scanners and exploit frameworks (e.g., Metasploit, Nmap, Nikto, WPScan) but rely on rigid rules and shallow evidence fusion, limiting adaptability to dynamic attack scenarios [3, 4, 13, 14, 15]. RL-based agents formulate pentesting as a Partially Observable Markov Decision Process (POMDP) with reward shaping, providing a principled approach to automating attack strategies [5, 6, 16]. Notably, the Raiju framework has made significant strides in automating post-exploitation tasks by leveraging RL algorithms, specifically Advantage Actor-Critic (A2C) and Proximal Policy Optimization (PPO) [17]. Integrated with Metasploit, Raiju trains specialized agents to perform tasks such as privilege escalation, hashdump gathering, and lateral movement in real-world environments, achieving a success rate exceeding 84% across diverse attack types in four tested environments. However, RL-based approaches, including Raiju, face two primary challenges: (i) the need for extensive state and action space engineering, and (ii) limited cross-target transferability without costly retraining. These limitations highlight the need for more adaptive methods, such as LLMs, to enhance efficiency and generalization in penetration testing.
2.2 LLM single-/few-agent pipelines
@@ 2.3 LLM multi-agent orchestration
PentestAgent adopts RAG-grounded, role-based collaboration (reconnaissance, triage, exploitation) to reduce hallucinations and improve next-step selection [8]. Additionally, VulnBot structures collaboration via a Penetration Task Graph (PTG) to preserve phase order (recon $\rightarrow$ scanning $\rightarrow$ exploitation) and constrain branching; reported results include 30.3% overall and 69.05% subtask completion on AutoPenBench and strong performance on AI-Pentest-Benchmark [9, 10, 11]. RefPentester [19] introduces knowledge-informed self-reflection tied to stage recognition, improving recovery from failed operations on Hack The Box targets. RapidPen targets the initial foothold (IP-to-shell) with a ReAct-style loop and retrieval of exploit knowledge, demonstrating fully autonomous compromises on HTB within minutes at modest cost [20]. The work of Weber et al. [21] presents Perses, a notable effort to enable small language models (SLMs) to perform automated privilege escalation through an extensible, role-specialized multi-LLM architecture. Perses shows that heterogeneity—assigning lightweight models to Planner, Commander, Summariser and domain-specific Overseers—can substantially improve exploitation of simple misconfigurations. Importantly, the evaluation in Perses is narrowly scoped: experiments are conducted primarily on FreeBSD targets, employ a limited and largely handcrafted set of privilege-escalation vulnerabilities, and use a threat model tailored to configuration errors rather than broad end-to-end attacks. As a result, Perses demonstrates the viability of SLM heterogeneity in constrained environments but leaves open questions about transferability to full penetration pipelines (reconnaissance, scanning, multi-stage exploitation), complex real-world services, and heterogeneous network topologies.
@@
Fang et al. [22] show that, given CVE descriptions, GPT-4 can exploit 87% of a 15 one-day vulnerability set, whereas other LLMs and scanners achieve 0%; without the description, success drops markedly (7%). Zhu et al. [23] extend to teams of agents (HPTSA) for zero-day web vulnerabilities, reporting up to 42% pass@5 and 18% pass@1 on 14 real-world cases with GPT-4. For broader skill evaluation, HackSynth proposes a two-module agent and two CTF-based benchmarks (PicoCTF/OverTheWire; 200 tasks) [24], while NYU CTF Bench (NeurIPS D&B) contributes a scalable open-source dataset and automation framework (200 CSAW CTF tasks) [25].
@@ 2.5 Benchmarks and methodology
The emergence of AI-driven penetration testing has been accompanied by a rapid proliferation of evaluation suites designed to measure autonomy, tool integration, and end-to-end performance under controlled conditions. While the space remains nascent, a few benchmarks have begun to dominate experimental protocols. Notably, AutoPenBench and AI-Pentest-Benchmark appear most frequently in recent studies, reflecting their alignment with realistic, multi-phase pentesting workflows and their ability to grade performance across autonomy levels and subtasks. Conversely, more specialized testbeds such as CVE-Bench target specific exploitability dimensions (e.g., real-world CVEs in web contexts) and thus see adoption in works focusing on vulnerability exploitation rather than full-cycle orchestration. Capture-the-flag (CTF)-oriented resources such as NYU CTF Bench and the datasets introduced by HackSynth have also gained traction, particularly for skill-granular or task-decomposed evaluations, though their scenarios often differ from operational pentests in scope and realism.
Within this landscape, AutoPenBench offers open, graded tasks spanning web, network, and cryptographic targets, with configurable autonomy modes to support comparisons between orchestration strategies and model backbones [10]. AI-Pentest-Benchmark provides VM-based targets, enabling reproducible end-to-end penetration tests [11], thereby supporting performance attribution across discovery, exploitation, and post-exploitation phases. CVE-Bench grounds evaluation in real-world web CVEs, reporting typical success rates in the low teens even for state-of-the-art agents, highlighting the gap between research prototypes and robust autonomy [26]. Methodological recommendations across these works increasingly emphasize standardizing budget constraints, clearly labeling autonomy levels, and reporting detailed error modes to prevent overestimation of capabilities [27].

### v2

# xOffense: An Autonomous Multi-Agent Framework for Penetration Testing with Domain-Adapted Large Language Models

@@ Abstract
Penetration testing plays a critical role in assessing the security of modern information systems; however, existing automated approaches based on machine learning, deep learning, or reinforcement learning remain constrained by simplified action spaces, high computational overhead, and limited reasoning across multi-stage workflows such as reconnaissance, vulnerability analysis, and exploitation. Recent large language model (LLM)-based systems have shown promise in addressing these challenges, yet they often rely on large-scale or proprietary models, resulting in high cost, limited scalability, and suboptimal adaptability in complex environments. This paper introduces xOffense, an AI-driven multi-agent framework for autonomous penetration testing that transforms traditional expert-driven processes into fully automated and scalable workflows. The proposed system leverages a fine-tuned mid-scale open-source LLM to perform structured reasoning and decision-making, while decomposing the pentesting pipeline into specialized agents responsible for reconnaissance, vulnerability scanning, and exploitation. An orchestration mechanism coordinates inter-agent collaboration to ensure coherent multi-phase execution. In addition, domain-specific fine-tuning with chain-of-thought penetration testing data enables accurate command generation and consistent multi-step reasoning across tasks. The effectiveness of xOffense is validated on two representative benchmarks, AutoPenBench and AI-Pentest-Benchmark. Experimental results demonstrate that the proposed framework consistently outperforms existing LLM-based approaches, achieving a sub-task completion rate of 79.17% and surpassing state-of-the-art systems in both effectiveness and reliability. These results highlight that integrating domain-adapted mid-scale LLMs within a structured multi-agent architecture can provide a cost-efficient, scalable, and reproducible solution for autonomous penetration testing, offering a practical direction for deploying intelligent cybersecurity systems in real-world settings.
@@
Autonomous penetration testing , Large language models , Multi-agent systems , Intelligent decision-making , Cybersecurity automation , Domain-adapted learning
@@ 1 Introduction
Pentest remains one of the most effective ways to assess the real-world security posture of modern information systems. Unfortunately, the prevailing approach manual testing conducted by small teams of human experts cannot keep pace with today’s rapidly expanding attack surface. In 2024 alone, the National Vulnerability Database listed more than 29,000 new CVEs, a 38% year-over-year increase [21] [8]. As networks grow in scale and complexity, the gap between the appearance of new vulnerabilities and the ability of security professionals to detect and remediate them is widening. This growing imbalance underscores the urgent necessity for automated and intelligent pentest solutions.
@@
A second, more adaptive line of work leverages RL. Systems like IAPTS [32] and HA-DRL [39] model pentest as a sequential decision-making problem in partially observable environments, enabling agents to autonomously explore and learn effective attack strategies through interaction and reward-based learning. RL agents can, in principle, discover novel attack paths, yet in practice they face two key obstacles: (i) their action space must be heavily simplified such as “scan port”, “exploit CVE-xxx”, and (ii) training requires a large number of environment interactions that are expensive to obtain and seldom transfer between real networks. Consequently, even state-of-the-art RL-based pentesters achieve modest coverage and require significant engineering to integrate new tools or protocols.
@@
These limitations illustrate that purely DL or RL-based automation is insufficient for the inherently multi-phase and dynamic nature of pentest. To overcome this, recent works have turned toward AI agent-based paradigms, in which multiple specialized agents collaborate to emulate the workflow of human red teams. In such systems, each agent assumes a distinct role: a Reconnaissance Agent focuses on host and service discovery, a Vulnerability Analysis Agent correlates findings with CVE and CWE knowledge bases, and an Exploitation Agent generates and tests candidate payloads. This agent-oriented decomposition enables modularity, context retention across phases, and the possibility of scaling to complex attack paths that traditional ML/DL/RL pipelines cannot handle.
@@
Recent advancements in LLMs have opened new possibilities for automating modern pentest. Leveraging their strong reasoning and code generation capabilities, LLMs have been adopted in several research prototypes such as PentestGPT [5], PentestAgent [34], and VulnBot [17], where models assist or autonomously conduct reconnaissance, scanning, and exploitation. In particular, VulnBot represents a major step forward: it frames pentest as a collaborative workflow between specialized LLM agents guided by a Penetration Task Graph (PTG), enabling the simulation of expert-level pentesting with limited or no human intervention. Empirical results from benchmarks like AutoPenBench [9] and AI-Pentest-Benchmark [16] have validated VulnBot’s capacity to outperform other automated methods in structured testing environments.
@@
However, most of these systems rely heavily on extremely large or commercial LLMs, such as GPT-4o, LLaMA3-70B, or DeepSeek-V3. Despite their capabilities, these models present significant operational hurdles including high resource consumption, costly API dependencies, and limited adaptability to domain-specific fine-tuning. Additionally, their general-purpose nature often leads to hallucinations, loss of context across phases, or poor command translation in complex penetration workflows. As such, there is a pressing need to explore whether smaller, fine-tuned open-source models can serve as more efficient, specialized alternatives, which offers better controllability, lower cost, and targeted reasoning.
@@
This work is motivated by two core observations. First, while large LLMs have demonstrated strong potential in security domains, scale alone does not guarantee effectiveness, particularly when models are deployed in structured, multi-step tasks like pentest. Despite their size, large-scale LLMs still suffer from context loss across phases, generate incorrect tool usage, and require significant human supervision. Second, most current systems adopt LLMs as black-box assistants or instruction followers, without integrating deeper task-specific guidance or domain adaptation. To address these limitations, we explore an alternative paradigm: leveraging a mid-sized, domain-adapted LLM that is explicitly trained for pentest tasks. We present xOffense, a refined evolution of the VulnBot framework that substitutes its dependence on large, general-purpose models with a fine-tuned Qwen3-32B [45], a 32-billion-parameter open-source language model. Through dedicated training on pentest workflows, including vulnerability scanning, exploit crafting, and security tool interaction, xOffense achieves sharper task alignment, enhanced operational fidelity, and greater adaptability in nuanced or low-visibility environments.
@@
Beyond simply replacing the core language model, xOffense also incorporates a context-aware prompting scheme we refer to as grey-box prompting. In this setup, agents are equipped with partial system insights, such as protocol hints, observed services, or prior scan summaries, enabling them to make more informed decisions without relying on full system disclosure. This strategy preserves the operational constraints of black-box testing while offering minimal structured guidance, striking a balance between realism and agent effectiveness. By preserving VulnBot’s three-phase pipeline reconnaissance, scanning, and exploitation xOffense ensures compatibility with existing workflows, facilitates direct benchmarking, and provides a robust foundation for comparative evaluation.
@@
In this paper, we present the design, implementation, and evaluation of xOffense, a lightweight, domain-adaptive, and highly effective autonomous pentest system. Our key contributions are as follows:
@@
    An AI-driven multi-agent pentest system. We propose a novel agent-based framework in which specialized agents collaborate to cover all critical phases of pentest reconnaissance, vulnerability analysis, and exploitation. This design emulates the workflow of human red teams, ensures modularity across tasks, and enables coherent orchestration of complex attack paths in an autonomous manner.
@@
    A domain-adapted mid-scale LLM. At the core of our system lies Qwen3-32B, a 32B-parameter open-source model fine-tuned with Chain-of-Thought (CoT) pentest data. This adaptation empowers the model with precise multi-phase reasoning, accurate tool command generation, and strong adaptability in complex exploitation workflows.
@@
    Grey-box phase prompting. We introduce a context-aware prompting mechanism that selectively integrates environmental cues such as observed protocols, discovered services, and prior scan outputs into the agent reasoning process. This strategy strikes a balance between black-box and white-box testing, reducing context loss and improving continuity across phases.
@@
    Extensive empirical validation. We conduct rigorous evaluations of xOffense on AutoPenBench and AI-Pentest-Benchmark, demonstrating state-of-the-art performance in both synthetic and real-world pentest scenarios. The system achieves superior task and sub-task completion rates compared to prior methods, confirming the effectiveness of multi-agent orchestration and domain-adapted LLMs.
The remainder of this paper is organized as follows. Section 2 reviews prior studies on pentest and automation approaches, while Section 3 introduces the fundamental concepts that underpin automated systems. Section 4 describes the architecture of the proposed xOffense framework, including its fine-tuned Qwen3-32B model, grey-box prompting strategy, and multi-agent orchestration. We describe the experimental settings and benchmark datasets, evaluation metrics in Section 5. In Section 6 empirical results on two benchmarks and real-world exploitation scenarios are reported and analyzed. Section 7 discusses potential threats to validity and their implications for generalizability. In Section 8, we discuss ethical considerations and responsible use of the proposed framework, including potential misuse risks, deployment constraints, and implications for defensive cybersecurity research. Finally, Section 9 concludes the paper and outlines directions for future research.
@@ 2.1 Pre-LLM automation and RL
Deterministic orchestrators, such as DeepExploit, integrate scanners and exploit frameworks, including Metasploit, Nmap, Nikto, and WPScan, but rely on rigid rules and shallow evidence fusion, limiting adaptability to dynamic attack scenarios [36, 31, 11, 2, 43]. RL-based agents formulate pentesting as a Partially Observable Markov Decision Process (POMDP) with reward shaping, providing a principled approach to automating attack strategies [32, 39, 20]. Notably, the Raiju framework has made significant strides in automating post-exploitation tasks by leveraging RL algorithms, specifically Advantage Actor-Critic (A2C) and Proximal Policy Optimization (PPO) [28]. Integrated with Metasploit, Raiju trains specialized agents to perform tasks such as privilege escalation, hashdump gathering, and lateral movement in real-world environments, achieving a success rate exceeding 84% across diverse attack types in four tested environments. However, RL-based approaches, including Raiju, face two primary challenges: (i) the need for extensive state and action space engineering, and (ii) limited cross-target transferability without costly retraining. These limitations highlight the need for more adaptive methods, such as LLMs, to enhance efficiency and generalization in pentest.
2.2 LLM single-few-agent pipelines
@@ 2.3 LLM multi-agent orchestration
PentestAgent adopts RAG-grounded, role-based collaboration (reconnaissance, triage, exploitation) to reduce hallucinations and improve next step selection [34]. Additionally, VulnBot structures collaboration via a PTG to preserve phase order (recon $\rightarrow$ scanning $\rightarrow$ exploitation) and constrain branching; reported results include 30.3% overall and 69.05% sub-task completion on AutoPenBench and strong performance on AI-Pentest-Benchmark [17, 9, 16]. RefPentester [3] introduces knowledge-informed self-reflection tied to stage recognition, improving recovery from failed operations on Hack The Box targets. RapidPen targets the initial foothold (IP-to-shell) with a ReAct-style loop and retrieval of exploit knowledge, demonstrating fully autonomous compromises on HTB within minutes at modest cost [23]. The work of Weber et al. [42] presents Perses, a notable effort to enable small language models (SLMs) to perform automated privilege escalation through an extensible, role-specialized multi-LLM architecture. Perses shows that heterogeneity, which assigns lightweight models to Planner, Commander, Summariser and domain-specific Overseers, can substantially improve exploitation of simple misconfigurations. Importantly, the evaluation in Perses is narrowly scoped: experiments are conducted primarily on FreeBSD targets, employ a limited and largely handcrafted set of privilege escalation vulnerabilities, and use a threat model tailored to configuration errors rather than broad end-to-end attacks. As a result, Perses demonstrates the viability of SLM heterogeneity in constrained environments but leaves open questions about transferability to full penetration pipelines (reconnaissance, scanning, multi-stage exploitation), complex real-world services, and heterogeneous network topologies.
@@
Fang et al. [7] show that, given CVE descriptions, GPT-4 can exploit 87% of a 15 one-day vulnerability set, whereas other LLMs and scanners achieve 0%; without the description, success drops markedly (7%). Zhu et al. [48] further extend this to teams of agents (HPTSA) for zero-day web vulnerabilities, reporting up to 42% pass@5 and 18% pass@1 on 14 real-world cases with GPT-4. For broader skill evaluation, HackSynth proposes a two-module agent and two CTF-based benchmarks (PicoCTF/OverTheWire; 200 tasks) [22], while NYU CTF Bench (NeurIPS D&B) contributes a scalable open-source dataset and automation framework (200 CSAW CTF tasks) [33].
@@ 2.5 Benchmarks and methodology
The emergence of AI-driven pentest has been accompanied by a rapid proliferation of evaluation suites designed to measure autonomy, tool integration, and end-to-end performance under controlled conditions. While the space remains nascent, a few benchmarks have begun to dominate experimental protocols. Notably, AutoPenBench and AI-Pentest-Benchmark appear most frequently in recent studies, reflecting their alignment with realistic, multi-phase pentesting workflows and their ability to grade performance across autonomy levels and subtasks. Conversely, more specialized testbeds such as CVE-Bench target specific exploitability dimensions, such as real-world CVEs in web contexts, and thus see adoption in works focusing on vulnerability exploitation rather than full-cycle orchestration. Capture The Flag resources such as NYU CTF Bench and the datasets introduced by HackSynth have also gained traction, particularly for skill-granular or task-decomposed evaluations, though their scenarios often differ from operational pentests in scope and realism.
Within this landscape, AutoPenBench offers open and standardized graded tasks spanning web, network, and cryptographic targets, with configurable autonomy modes to support comparisons between orchestration strategies and model backbones [9]. AI-Pentest-Benchmark provides VM-based targets, enabling reproducible end-to-end penetration tests [16], thereby supporting performance attribution across discovery, exploitation, and post-exploitation phases. CVE-Bench grounds evaluation in real-world web CVEs, reporting typical success rates in the low teens even for state-of-the-art agents, highlighting the gap between research prototypes and robust autonomy [47]. Methodological recommendations across these works increasingly emphasize standardizing budget constraints, clearly labeling autonomy levels, and reporting detailed error modes to prevent overestimation of capabilities [15].